Find the domain and range of the function
f(x)=1/2-sin3x
1
Expert's answer
2020-10-12T13:50:55-0400

We can calculate value "f(x)=\frac{1}{2}-\sin{3x}" for every value of "x", so domain of function is "(-\infty, +\infty)". The range of "(-\sin{3x})" is "[-1, 1]", and if we add "\frac{1}{2}" ,

range of "f(x)" is "[-\frac{1}{2}, 1\frac{1}{2}]" .